# USTC-TD: A Test Dataset and Benchmark for Image and Video Coding in 2020s

### v2

@@ V-B Experimental Results
TABLE XI: BD-rate (%) Comparison for PSNR. The anchor is VTM.
Dataset VTM HM DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
@@
USTC-TD 0.0 46.9 286.7 133.0 69.3 69.8 63.9 17.0 9.5 4.5 7.1 23.1
Average 0.0 41.1 229.6 139.4 71.3 44.5 36.3 1.5 -9.6 -11.1 -16.3 -10.2
TABLE XII: BD-rate (%) Comparison for MS-SSIM. The anchor is VTM.
Dataset VTM HM DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
@@
USTC-TD 0.0 47.9 38.6 -18.1 77.5 -41.1 -62.1 -62.3 8.4 -66.3 -68.6 37.1
Average 0.0 39.4 71.9 28.8 41.9 -23.2 -30.3 -47.9 -11.3 -50.3 -55.1 -10.7
V-B1 Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD Image Datasets
Taking bpp as the horizontal axis and the reconstructed PSNR/MS-SSIM as the vertical axis, we present the rate and distortion curves of different image compression schemes over USTC-TD 2022 and 2023 image datasets in Fig. 6 . From the overall results, we can find that the partial learned schemes (iWave++[9], ELIC[10], MLIC++[11]) can outperform the traditional image compression schemes and achieve better compression performance than H.266/VVC on proposed datasets under different metrics, which show its powerful potential. Here we further analyze their performance from the perspective of content factors of different test images of USTC-TD.
As shown in Fig. 8 , the detailed RD curves of some test images (USTC-2022-12, USTC-2023-15, USTC-2022-17, USTC-2023-07) with some special phenomenons are illustrated, and the results of each test image are presented in the Section II of supplementary material. Based on the performance comparison of these schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include the following four aspects:
  • (4) Compared to the performance of these schemes on different datasets[28, 29, 30], the proposed datasets can collaborate with other datasets to handle a wide coverage of performance evaluation, which also demonstrates the efficiency of specific design of the proposed datasets’ different factors/features.
@@ V-B1 Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD Image Datasets
V-B2 Evaluation and Analysis of Advanced Video Compression Schemes on USTC-TD Video Dataset
@@
Taking bpp as the horizontal axis and the reconstructed PSNR/MS-SSIM as the vertical axis, we present the rate and distortion curves and BD-rate results of different video compression schemes over the USTC-TD 2023 video dataset in Fig. 7 , Table XI  and Table XII . From the overall results of PSNR metric, we can find that the performance of the advanced traditional schemes is better than that of the all advanced learned schemes on the proposed dataset, which is opposed to the performance on other datasets [31, 32, 33, 34] . From the overall results of MS-SSIM metric, the performance of the traditional schemes is lower than that of advanced learned schemes. Here we further analyze their performance from the perspective of typical characteristics of different test video contents of USTC-TD video dataset.
As shown in Fig. 9 , the detailed RD curves of some test videos (USTC-FourPeople, USTC-BasketballPass, USTC-Snooker, USTC-BicycleDriving) with some special phenomenons are illustrated, the results of each test video are also presented in the Section II of supplementary material. Based on the performance comparison of advanced schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include the following three aspects:
@@
Figure 11: Visualization of limitation of partial advanced video compression schemes (TCM-VC[20], DCVC-DC[26], DCVC-FM[27]) on USTC-BicycleDriving and USTC-Snooker.
  • (1) The traditional schemes show the robust generalization ability on various real-world scenarios. As shown in Table XI , different from the performance of advanced compression schemes on other datasets [31, 32, 33, 34] , the traditional schemes still achieve the state-of-the-art performance of these schemes on USTC-TD. Different from the other datasets, USTC-TD video dataset focuses on various temporal correlation types. Combined the attributes of different test videos of USTC-TD (Fig. 5 , Table VIII ), we find that the traditional schemes can handle more scenarios with the various temporal features of motion-related elements (motion type, lens motion), such as the USTC-BicycleDriving with higher scores of TI and USTC-Snooker with specific design of lens motion (mentioned in Section IV.C). The performance of these test videos is shown in Fig. 9  (c) and (d), and the results demonstrate that the traditional schemes can handle these scenes with complex motion types robustly.
  • (2) The learned schemes show the optimistic potential on some scenarios with complex motion, such as the USTC-FourPeople with lower scores of TI, and the USTC-BasketballPass with higher scores of TI. These scenarios commonly appear in the previous test datasets[31, 32, 33, 34] , such as the FourPeople and BasketballPass in HEVC/VVC CTC, which can guide the design and the optimized target of deep network to handle these motion situations. The results of these specific scenarios demonstrate the basic evaluative ability of the previous datasets and the performance potential of the deep learning-based manner.
  • (3) Based on the analysis of video-related features (Section IV.C), our proposed dataset can make up more typical motion/temporal correlation-related real-world factors with other datasets to handle a wide coverage of performance evaluation, which also demonstrates the efficiency of the specific design of the proposed datasets’ different factors/features. Furthermore, compared to the performance of other datasets[31, 32, 33, 34] of different schemes mentioned in [7, 8, 9, 10, 11], as shown in Table XI  and Table XII , the different phenomenon between the performance of our proposed dataset and other datasets also verifies the efficiency of the proposed dataset.
V-C Limitation and Inspiration
In this subsection, based on the analysis of experimental results, we further analyze the limitations of evaluative image/video compression schemes, and point out some limitations and inspirations among these advanced compression schemes.
@@ V-C1 Limitation and Inspiration of Image Compression
Based on the above analysis in Section V.B, we can find that the generalization ability of the learned codec is a major challenge for practical usage. Most existing methods focus on improving the compression performance while neglecting its generalization for various scenarios. As the situations mentioned in the item (1) and (2) of conclusions (Section V.B (1)), the generalization ability is challenged with the effective extension of the evaluation data. These problems mainly arise from the incompletion of training data and the constraint optimization direction of deep learning-based manner with the limited evaluation data. Here we further explore these problems based on the coding process of different compression schemes.
@@ V-C2 Limitation and Inspiration of Video Compression
Based on the above analysis in Section V.B, as shown in Fig. 7 and Table XI , the performance of all learned video codecs is lower than that of traditional video codecs on USTC-TD, which is different from the performance phenomenon on other datasets. The reason mainly comes from that the learned video codecs perform poorly on some sequences with complex motion features. Here we further explore these problems based on the motion-related modules of different schemes.
@@
TABLE XIII: BD-rate (%) Results of Different Training Strategies of Optical Flow-related Module of DCVC-DC
on USTC-TD [26] under PSNR Metric
      Scheme       BD rate (%)
      DCVC-DC       7.1%
      DCVC-DC + Flow Pre-training       -1.6%
      $\triangle$       -8.7%
@@
To further verify the performance impact of these problems, we tentatively design the experiment to optimize the optical flow-related module of different learned video codecs. We set the state-of-the-art scheme (DCVC-DC [26]) as the anchor, and use the motion vectors of H.266/VVC as the optimized target of the optical flow-based motion estimation module (Spynet[51]) in the offline pre-training stage, instead of the usage of EPE loss for the training of these optical flow-based modules. The performance is shown in Table XIII . Inspired by the results, it verifies that the motion modeling and training strategy of learned video compression models are necessary to be further improved for practical usage in the future.
VI Conclusion

### v3

@@ V-B Experimental Results
V-B1 Objective Quality Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD
Taking bpp as the horizontal axis and the reconstructed PSNR, MS-SSIM, VMAF (MSE/MS-SSIM model) as the vertical axis, we present the rate and distortion curves of different image compression schemes over USTC-TD 2022 and 2023 image datasets in Fig. 6 . From the overall results, we can find that the partially learned schemes (iWave++[9], ELIC[10], MLIC++[11]) can outperform the traditional image compression schemes and achieve better compression performance than H.266/VVC on proposed datasets under different metrics, which show its powerful potential. Here we analyze their performance from the perspective of content factors of different test images of USTC-TD. As shown in Fig. 9 , the detailed RD curves of some test images (USTC-2022-12, USTC-2023-15, USTC-2022-17, USTC-2023-07) with some special phenomenons are illustrated, and the results of each test image are presented in supplementary material. Based on the performance comparison of these schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include four aspects:
  • (4) Compared to the performance of these schemes on different image datasets[32, 33, 34], the proposed datasets can collaborate with other datasets to handle a wide coverage of performance evaluation, which also demonstrates the efficiency of the specific design of the proposed datasets’ different content factors/features.
(a) USTC-2022-12
(b) USTC-2023-15
(c) USTC-2022-17
(d) USTC-2023-07
(a) USTC-FourPeople
(b) USTC-BasketballPass
(c) USTC-BicycleDriving
(d) USTC-Snooker
Figure 9: Specific rate-distortion (RD) curves of advanced video compression schemes on partial evaluative short videos under PSNR metric.
TABLE XII: The Overall MOS Results of Compressed Images of Classic Standardized and Advanced Learned Image Compression Schemes, where Blue Represents the Lowest Score and Red Represents the Highest Score.
Dataset
Scheme
Classification
Compression
Scheme
MOS
Score
USTC-TD 2022 & 2023 Image Dataset Traditional H.266/VVC [2] 3.67
Learned Factorized Model[6] 3.44
Hyperprior Model[6] 3.43
Autoregressive Model[7] 3.65
Cheng2020[8] 3.77
MLIC++ [11] 3.80
V-B2 Subjective Quality Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD
In Table XII , we present the overall MOS results of compressed images of different image compression schemes over USTC-TD 2022 and 2023 image datasets. From the overall results, we can find that recent advanced learned compression schemes [8, 11] outperform the traditional compression schemes, which demonstrates their powerful potential. This can be attributed to the optimization of MS-SSIM for perceptual quality and the fact that end-to-end learning methods, compared to hand-crafted approaches, are more flexible and can be easily optimized toward different quality assessment metrics. From the detailed results of compressed images of different schemes mentioned in supplementary material, it is observed that different compression schemes show different MOS performance trends on different images. Here we analyze their performance from the perspective of feature analysis of different test images of USTC-TD (Section IV.C), and explore the advantages of different learned compression schemes for varying scenes under the evaluation of subjective metric. The conclusions include the following two aspects:
  • (1) The learned schemes show the powerful potential on various scenarios for perception optimization, especially the complex scenarios with mixture features, such as the USTC-2023-03 with the lower SI and higher CF scores, USTC-2023-05 with the higher SI, lower CF and LI scores. Combined with the performance of objective quality metrics, these samples highlight the learned schemes’ generalization in optimizing across different metrics.
  • (2) Despite traditional schemes[2] being hand-crafted and optimized for PSNR, their subjective optimization capability remains competitive with certain learned schemes. Moreover, they exhibit a degree of practical applicability for visual perception, achieving comparable performance to some state-of-the-art learned schemes [8, 11] in challenging scenarios, such as USTC-2022-02 with lower SI and LI scores, and USTC-2023-15 with higher SI and lower LI scores. Therefore, for the development of learned schemes, merely modifying the optimization metric is not sufficient; it is essential to further integrate perceptual theories into the framework design to achieve more effective improvements.
TABLE XIII: BD-rate (%) Comparison for PSNR. Short Setting with Intra Period = 32. The anchor is VTM.
Dataset VTM HM AV1 AV2 DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
@@ V-B2 Subjective Quality Evaluation and Analysis of Advanced Image Compression Schemes on USTC-TD
USTC-TD1 0.0 46.7 34.3 11.0 284.4 124.7 64.6 67.1 60.4 16.0 8.2 3.7 6.3 24.9
Average (Desirable) 0.0 41.8 230.9 139.7 72.8 47.2 38.4 4.0 -8.0 -10.0 -15.1 -8.4
Average (Ideal) 0.0 41.1 229.3 138.4 70.7 44.2 35.8 1.4 -9.8 -11.2 -16.4 -9.9
  • 1
    The results of USTC-TD indicate the average results of the Challenging preset of dataset collaboration mentioned in Table XI  .
TABLE XIV: BD-rate (%) Comparison for MS-SSIM. Short Setting with Intra Period = 32. The anchor is VTM.
Dataset VTM HM AV1 AV2 DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
@@
USTC-TD1 0.0 47.7 31.8 11.0 55.8 3.9 3.4 -19.1 -23.7 -44.3 0.4 -48.0 -48.7 39.5
Average (Desirable) 0.0 39.6 74.0 34.3 32.5 -22.5 -27.1 -47.5 -12.5 -50.0 -54.8 -10.8
Average (Ideal) 0.0 39.3 74.1 31.5 32.6 -20.5 -25.5 -45.7 -12.3 -48.0 -52.6 -10.4
  • 1
    The results of USTC-TD indicate the average results of the Challenging preset of dataset collaboration mentioned in Table XI  .
TABLE XV: BD-rate (%) Comparison for VMAF (MSE model), VMAF (MS-SSIM model).
Short Setting with Intra Period = 32. The anchor is VTM.
Dataset Evaluative Metric VTM HM AV1 AV2 DVC_Pro DCVC CANF-VC TCM-VC VNVC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
USTC-TD VMAF (MSE model) 0.0 46.5 291.9 137.1 118.5 63.0 39.6 45.8 42.2 10.6 9.5 -1.1 9.2 21.6
VMAF (MS-SSIM model) 160.6 89.4 93.7 107.3 112.8 64.8 46.1 34.5
TABLE XVI: BD-rate (%) Comparison for PSNR, MS-SSIM, VMAF (MSE model), VMAF (MS-SSIM model).
Long Setting with Intra Period = 32. The anchor is VTM.
Dataset Evaluative Metric VTM AV1 AV2 DVC_Pro DCVC CANF-VC TCM-VC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
USTC-TD PSNR 0.0 33.6 11.0 403.5 120.8 70.2 54.4 9.5 2.8 -1.5 0.4 13.5
MS-SSIM 0.0 28.4 7.0 43.3 5.6 5.7 -25.3 -49.1 -3.0 -51.5 -53.4 23.5
VMAF (MSE model) 0.0 349.2 194.5 118.9 58.0 23.8 33.7 4.5 3.5 -6.7 2.4 10.5
VMAF (MS-SSIM model) 166.8 104.6 94.8 111.0 69.9 52.1 37.8
TABLE XVII: BD-rate (%) Comparison for PSNR, MS-SSIM, VMAF (MSE model), VMAF (MS-SSIM model).
Long Setting with Intra Period = -1. The anchor is VTM.
Dataset Evaluative Metric VTM AV1 AV2 DVC_Pro DCVC CANF-VC TCM-VC DCVC-HEM OOFE SDD DCVC-DC DCVC-FM
USTC-TD PSNR 0.0 33.4 -1.9 520.7 221.4 95.6 99.1 23.9 16.5 15.2 9.2 22.4
MS-SSIM 0.0 24.4 -5.0 64.9 30.7 17.5 -2.3 -41.3 10.7 -42.5 -51.2 34.7
VMAF (MSE model) 0.0 387.8 95.7 195.4 100.0 49.1 59.2 17.3 15.1 4.1 10.4 18.7
VMAF (MS-SSIM model) 277.2 132.7 146.3 158.0 99.0 82.9 50.8
@@
V-B3 Objective Quality Evaluation and Analysis of Advanced Video Compression Schemes on USTC-TD
@@
Taking bpp as the horizontal axis and the reconstructed PSNR, MS-SSIM, VMAF (MSE/MS-SSIM model) as the vertical axis, we present the rate and distortion curves and BD-rate results of different video compression schemes over the USTC-TD 2023 video dataset in Fig. 7 , Table XVII, XVII , XVII  for short dataset, Table XVII , Table XVII  for long dataset. Note that the setting of intra period = -1 is used to enhance comprehensive benchmark for the testing of long sequences, it allows for an accurate evaluation of the video compression scheme’s robustness to handle the long video sequences. From the overall results of PSNR and VMAF (MSE model) metric, we can find that the performance of the advanced traditional schemes is better than that of all advanced learned schemes on the proposed short/long video dataset. Notably, the PSNR results of short setting differ from the conclusions drawn from other datasets [35, 36, 37, 38]. From the overall results of MS-SSIM metric, the performance of the traditional schemes is lower than that of advanced learned schemes. Here we analyze their performance from the perspective of typical characteristics of different test video contents and test length of USTC-TD video dataset. As shown in Fig. 9 , the detailed RD curves of some test videos (USTC-FourPeople, USTC-BasketballPass, USTC-Snooker, USTC-BicycleDriving) with some special phenomenons are illustrated, the results of each test video are also presented in supplementary material. Based on the performance comparison of advanced schemes and feature analysis of proposed datasets (Section IV.C), the conclusions mainly include the following four aspects:
  • (1) The traditional schemes show good generalization ability on various real-world video scenarios. As shown in Table XVII , different from the performance of advanced compression schemes on other datasets [35, 36, 37, 38] , the traditional schemes still achieve the state-of-the-art performance of these schemes on USTC-TD under the PSNR metric. Different from the other datasets, USTC-TD video dataset focuses on various temporal correlation types. Combined the attributes of different test videos of USTC-TD (Fig. 5 , Table VIII ), we find that the traditional schemes can handle more scenarios with the various temporal features of motion-related elements (motion type, lens motion), such as the USTC-BicycleDriving with higher scores of TI and USTC-Snooker with the specific design of lens motion (Section IV.C). The performance of these test videos is shown in Fig. 9  (c) and (d), and the results demonstrate that the traditional schemes can handle these scenes with complex motion types robustly.
  • (2) The learned schemes show the optimistic potential on some scenarios with complex motion cases, such as the USTC-FourPeople with lower TI scores, and the USTC-BasketballPass with higher TI scores. These scenarios commonly appear in the previous test datasets[35, 36, 37, 38] , such as the FourPeople and BasketballPass in HEVC/VVC CTC, which can guide the design and the optimized target of deep network to handle these motion situations. The results of these specific scenarios demonstrate the basic evaluative ability of the previous datasets and the performance potential of the deep learning-based manner.
  • (3) Based on the performance comparison of advanced compression schemes under different test length settings and intra period settings of USTC-TD (Table XVII for short dataset, Table XVII  and Table XVII  for long dataset), the performance gap between traditional and learned schemes is amplified in the long setting, it further demonstrates the superiority of traditional schemes. Combined with the attributes of different length settings (Fig. 5 ), the more powerful temporal diversity of long sequences challenges the robustness of these schemes even further. The extended temporal variation with different intra period tests further exposes the shortcomings of advanced compression schemes, revealing the limitations of learned schemes in the error propagation of prediction chain, motion modeling, and the design of reference mechanism.
  • (4) Based on the analysis of video-related features (Section IV.C), our proposed dataset can make up more typical motion/temporal correlation-related real-world factors with other datasets to handle an excellent coverage of performance evaluation, which also demonstrates the efficiency of the specific design of the proposed datasets’ different factors/features. Furthermore, compared to the performance of other datasets[35, 36, 37, 38] of different schemes mentioned in [12, 16, 19, 20, 21, 22, 23, 25, 26, 27], as shown in Table XVII  and Table XVII , the different phenomenon between the performance of our proposed dataset and other datasets also verifies the efficiency of the proposed dataset.
  • (5) Based on the performance comparison of different presets of dataset collaborations under different feature coverage settings, as shown in Table XVII and XVII , the performance trends observed across different presets highlight their distinct testing preferences, and further validate the configuration rationality of different presets. Combining the attributes of each preset, the results from the Challenging preset to the other presets gradually validate the performance of these video compression schemes, highlighting the testing effects with the consideration of general, reliable, and comprehensive evaluation, respectively. Specifically, in Table XVII and XVII , the performance of DVC_Pro and CANF-VC tend to show variational BD-rate changes under different presets, reflecting their limitations in handling specific considerations or coverage settings compared to other methods. On the other hand, the variants of DCVC show more consistent performance across different feature coverage settings, demonstrating that these methods are more robust when adapting to various scenarios.
V-B4 Subjective Quality Evaluation and Analysis of Advanced Video Compression Schemes on USTC-TD
In Table XVIII , we present the overall MOS results of compressed videos of different video compression schemes over USTC-TD 2023 video dataset. The results of video compression follow the same trend as image compression, where learned schemes outperform traditional schemes, highlighting their advantages in perceptual quality optimization and adaptability to different quality assessment metrics. From the detailed results mentioned in the supplemental material, we present the detailed MOS results of compressed videos of different video compression schemes over USTC-TD 2023 video dataset. It is also observed that different schemes show different MOS performance trends on different videos. Here we further analyze their performance from the perspective of feature analysis of different test videos of USTC-TD (Section IV.C), and explore the advantages of different learned compression schemes for varying scenes under the evaluation of subjective metric. The conclusions mainly include the following two aspects:
  • (1) The learned video schemes also demonstrate powerful potential for perceptual optimization, particularly in complex motion scenarios such as USTC-ShakingHands and USTC-BasketballPass, which exhibit higher TI scores. Combined with the performance of objective quality metrics, these samples further highlight the learned schemes’ ability to generalize and optimize across different metrics.
  • (2) Despite traditional schemes[2] being hand-crafted and optimized for PSNR, they exhibit strong competitiveness compared to learned schemes. Especially in certain sequences with extreme motions, they demonstrate a greater advantage over learned schemes, such as USTC-BicycleDriving with the highest TI scores, and USTC-BasketballDrill with higher SI and TI scores. Therefore, for the development of learned schemes, merely modifying the optimization metric is also not enough. It is crucial to further incorporate video-related factors into the framework design, such as temporal consistency and motion fidelity, to achieve more effective improvements.
V-C Limitation and Inspiration of Advanced Image and Video Compression Schemes
In this subsection, based on the analysis of experimental results, we analyze the limitations of evaluative image/video compression schemes, and point out some limitations and inspirations among these advanced compression schemes.
TABLE XVIII: The Overall MOS Results of Compressed Videos of Classic Standardized and Advanced Learned Video Compression Schemes, where Blue Represents the Lowest Score and Red Represents the Highest Score.
Dataset
Scheme
Classification
Compression
Scheme
MOS
Score
USTC-TD 2023 Video Dataset Traditional H.266/VVC[2] 3.49
Learned DCVC[19] 3.14
TCM-VC[20] 3.24
DCVC-HEM[21] 3.42
OOFE[22] 3.51
SDD[23] 3.41
DCVC-DC[26] 3.51
DCVC-FM[27] 3.55
@@ V-C Limitation and Inspiration of Advanced Image and Video Compression Schemes
Figure 11: Visualization of limitation of partial video compression schemes (TCM-VC[20], DCVC-DC[26], DCVC-FM[27]) on USTC-BicycleDriving/Snooker.
@@ V-C1 Limitation and Inspiration of Image Compression
Based on the above analysis in Section V.B, we can find that the generalization ability of the learned codec is a major challenge for practical usage. Most existing methods focus on improving the compression performance while neglecting its generalization for various scenarios. As the situations mentioned in the item (1) and (2) of conclusions (Section V.B (1)), the generalization ability is challenged with the effective extension of the evaluation data. These problems mainly arise from the incompletion of training data and the constraint optimization direction of deep learning-based manner with the limited evaluation data. Here we explore these problems based on the coding process of different compression schemes.
TABLE XIX: BD-rate (%) Results of Different Training Strategies of Optical Flow-related Module of DCVC-DC [26]
on USTC-TD under PSNR Metric
      Scheme       BD-rate (%)
      DCVC-DC       6.3%
      DCVC-DC + Flow Pre-training       -1.2%
@@ V-C2 Limitation and Inspiration of Video Compression
Based on the above analysis in Section V.B, as shown in Fig. 7 and Table XVII , the performance of all learned video codecs is lower than that of traditional video codecs on USTC-TD, which is different from the performance phenomenon on other datasets. The reason mainly comes from that the learned video codecs perform poorly on some sequences with complex motion features. Here we further explore these problems based on the motion-related modules of different schemes.
@@
To further verify the performance impact of these problems, we tentatively design the experiment to optimize the optical flow-related module of different learned video codecs. We set the state-of-the-art scheme (DCVC-DC [26]) as the anchor, and use the motion vectors of H.266/VVC as the optimized target of the optical flow-based motion estimation module (Spynet[55]) in the offline pre-training stage, instead of the usage of EPE loss for the training of these optical flow-based modules. The performance is shown in Table XIX . Inspired by the results, it verifies that the motion modeling and training strategy of learned video compression models are necessary to be further improved for practical usage in the future.
VI Limitation Discussion
Beyond the existing image/video datasets and USTC-TD, here we further discuss the limitations of these image/video datasets. With the collaboration of these datasets, the overall coverage of different image/video features has become more comprehensive, yet there are still some missing data types.
First, one notable missing type is computer-generated (CG) content, such as animations, cartoon content, and screen content. Unlike the camera-captured content of existing datasets, these special images/sequences differ significantly in visual characteristics and elements, such as vibrant color schemes, exaggerated motion, and mixed textures. For example, as shown in Fig. 5 , the particular sample of the existing datasets with the highest temporal diversity is the animated video of MCL-JCV [36], which is the only sample from other datasets, aside from USTC-TD, with TI scores higher than 40. Second, AI-generated content [69, 70] is another crucial missing data type, which is created using advanced generative models, such as GANS and diffusion models. These images/videos introduce unique challenges for compression models, as they often contain novel visual patterns, surreal features, or unusual motion patterns, which significantly differ from the content captured by human creators. The absence of these above non-camera-captured content types reveals a limitation of existing datasets, which mainly focus on natural images/videos captured by cameras. As synthetic media becomes more common in fields like entertainment, gaming, digital art, and AI-generated content, the demand for their compression also increases, creating a need for datasets that include CG and AI-generated media. Therefore, addressing these limitations in the future would enhance compression models’ ability to handle a broader range of visual content.
VII Conclusion